# Quantification of the Self-Excited Emotion Dynamics in Online Interactions

### v1

@@ Abstract
Emotions are essential for guiding human behavior, particularly in social interactions. In modern societies, a growing share of human interactions are taking place online which has been shown to amplify and distort the expression and perception of emotions. However, the entanglement across different emotions is not fully understood. We use a multivariate Hawkes self-excited point process to model and calibrate the temporal expressions of six basic emotions in YouTube live chats. This allows us to understand interdependencies among emotions, but also to disentangle the influence from the video content and social interactions with peers. Positive emotions are found to be more contagious, while negative emotions tend to leave a longer-lasting impression on users’ memories. Furthermore, we quantify the endogeneity of online emotion dynamics and find that peer interactions drive user emotional expressions 3-5 times more than passive content consumption. This underscores the powerful incentives of social interactions and the potential risk of emotional manipulation through the use of modern chatbots.
Introduction
Emotions play a significant role in motivating human behaviors [1]. Social interactions are the main trigger for most emotions, and these emotions are conveyed through interpersonal communication [2]. Emotions thus facilitate interactions by creating emotional synchronization [3], and hence play a significant role in shaping collective narratives and influence fundamental societal issues such as political movements, public health campaigns, social justice initiatives, cultural trends, and community resilience in times of crisis [4, 5, 6].
In recent decades, there has been a significant increase in the proportion of human interactions occurring online. With advances in AI and the proliferation of online bots, it is becoming increasingly difficult to distinguish between humans and automated systems in online interactions [7, 8, 9]. This has prompted an active exploration of the role of emotions in digital engagements and its consequences for society. It has been shown that characteristics of the online environment, such as anonymity and the lack of non-verbal cues, amplify the expression and perception of emotions [10]. Emotional content on social media is more likely to be shared [11, 12]. Online users have been shown to favor emotionality over argument quality in information diffusion [13]. In particular, moral emotions increase the diffusion of political content in social networks [14]. Posts containing out-group animosity are significantly more likely to be shared on Facebook and Twitter [15]. The effects of “echo chambers” and heterogeneity of opinions both contribute to polarization online [16]. Furthermore, the emotion of online reviews is a reliable predictor of commercial success in various industries from movies to restaurants [17].
Despite widespread attention to the role of emotions on online platforms, we identify three limitations in the literature. First, there is a lack of consensus on the emotional responses on social platforms [18, 19]. Features that encourage favorable self-representation and social validation have been shown to induce a positivity bias [20, 21, 22]. At the same time, online interactions are driven by a cognitive negativity bias as illustrated by the disproportional focus on negative information [23, 24, 25]. Second, the study of online emotion diffusion has typically focused on the contagion of a single emotion, whereas interactions across different emotions have been less explored. Further, empirical evidence on the contagion effects of positive and negative emotions has been mixed [18]. Positive emotions have been shown to have larger contagion effects on social media platforms than negative ones [12, 26, 27]. Other evidence points to contagion effects in similar magnitudes for positive and negative emotions [28]. On the contrary, negative emotions, specifically anger, appear to be more contagious across social networks [29]. Third, it is unclear what type of online engagement influences the emotional states of users in what proportion. For instance, it has been shown that live comments during live events display amplified emotional intensity relative to regular comments [30]. However, the degree to which emotion intensity is influenced by live videos and peer interactions remains an unresolved question. More generally, it is a fundamental challenge to disentangle the effect of exogenous (exo) and endogenous (endo) influences in complex system dynamics [31].
@@ Introduction
Figure 1: (a) Screenshot of a YouTube live video on the Theranos scandal involving Elizabeth Holmes with a live chat section on the right. The video content is captured by transcripts (subtitles), which we use to proxy exogenous emotional stimuli for the live chats. The live chat section displays timestamped messages from users reacting to the video content in real-time as the live video streams. We highlight an example of a transcript labeled as sad in blue, and a live chat message labeled as expressing anger in red. (b,c) We visualize the extraction of emotions from the video and live chat in the screenshot above. We plot the arrival of live chat messages that are labeled as sad, angry, indicated with diamond markers in blue, red. The shared x-axis shows the time in the video in units of minutes. The y-axis is unit-less. We label emotions non-exclusively. For instance, the sentence “Holmes is a victim of the fake news media” is labeled as both sad and angry. We assume these emotions are generated by the latent, inhomogeneous intensity defined by expression (1). (d) Time-varying component of the emotion sad in the video (transcript). Blue dots annotate the arrival of transcripts that are labeled as sad. The black dashed line shows our interpolation of the time-varying component.
We address these limitations by studying the co-evolution of 6 basic emotions in YouTube live chats. We represent emotions from the video content as exogenous triggers, and emotions from the live chats as endogenous ones. We then calibrate a multivariate self-excited conditional Hawkes point process [32, 33] to understand and disambiguate the influence of video content and prior live chats, as well as the interdependence across different states of emotions. By estimating the Hawkes model parameters across 397 videos spanning more than 780 hours of content, we find that peer social interactions contribute to user emotional expressions 3-5 times more than passive content consumption. In addition, the source of emotion intensity is attributed mainly to spontaneous user expressions rather than the video influence. This suggests that online emotions grow and spread more effectively through communication incentives and social interactions. Further, we find that positive emotions - especially joy - are more prone to contagion than negative ones. However, we identify a negativity bias in the temporal dynamics whereby users have a longer-lasting direct memory of negative emotions in online interactions. Our study provides practical implications for the design of social media and live-stream platforms. First, increasing user exposure to joyful content can diffuse positivity across social communities. Second, active communication channels - such as live chats or bots - can much more effectively engage users emotionally compared to passive ones - such as video advertisements.
Results
Figure 2: Visualizations of estimated parameters $\nu^{e,f}$, $\mu^{e}_{0}$, $\alpha^{e,f}$, and $\gamma^{e}$ of the intensity defined by expression (1) from the maximisation of the log-likelihood function (5). The parameters are fitted ten times, each time sampling 60% (238) of the total 397 videos at random. Values and error bars are then obtained as sample mean and standard deviations across all 10 fits, respectively. Entry $\alpha^{e,f}$ of the $\alpha$ matrix represents excitation from emotion $f$ to emotion $e$ in the live chat, while $\gamma^{e}$ represents the characteristic time-scale over which past emotions trigger new emotions $e$. The $\nu$ matrix demonstrates how emotions in the video trigger emotions in the live chat, while $\mu_{0}$ illustrates the spontaneous baseline intensity of spontaneous emotion expression.
Live Chat Data
@@ Results
We examine the dynamics of emotions in online discussions through a collection of YouTube Live videos. YouTube is a video-sharing platform catered to a broad range of users across different cultures. YouTube Live allows the audience to chat with one another while the video is being played in real-time. We refer to the live discussions as live chats and a single user live comment as a live chat message, or just message. Videos with the live chat replay function record timestamped live discussions and allow viewers to retrieve the live experience by simultaneously replaying the video content and live chat messages. Figure 1(a) shows an example of a YouTube live video (left) with a live chat section (right) where viewers post messages visible to the entire audience as the video streams. Our sample includes 92,412 live chat messages across 397 videos, adding up to 780 hours of content and covering 27 topics (see SI Appendix for more statistics). Beyond messages, we also collect data of the video transcripts - or subtitles - which capture the video content.
Emotions from these texts are categorized by the widely adopted 6 basic emotions model which includes joy, surprise, anger, disgust, fear, and sadness [34]. To this end, we fine-tune a Roberta transformer model that assigns emotion labels to each text [35]. The assignment is non-exclusive in the sense that each text can have multiple emotion labels (see Methods). An example is shown in Figure 1(b,c).
@@
Modeling Emotion Interactions with a Multivariate Hawkes self-excited Conditional Point Process
@@
Emotions are contagious [36]. We therefore propose to describe the arrival of $|\mathcal{E}|=6$ different emotions as $|\mathcal{E}|$-dimensional self- and cross-exciting Hawkes point processes [37]. Hawkes point processes have been used to study self-exciting phenomenons such as earthquakes [38], financial trades [39, 40], video viewing activities, [41], and information cascades [42]. The point process describes the statistical occurrence of different types of emotions, also called events. For a given emotion $e\in\mathcal{E}$, we distinguish two types of events: emotion $e$ is present in (the subscript of) the video content or emotion $e$ is present in the live chat. It is our goal to understand to what extent a given emotion $f\in\mathcal{E}$ triggers another emotion $e\in\mathcal{E}$ either exogenously from the video to the live chat or endogenously from live chats to live chats. Here the adjectives exogenous and endogenous refer to the point of view of the live chat. We assume that emotions in the live chat cannot influence emotions in the video. This is precisely the case if the video is pre-recorded (e.g. a documentary, as in Figure 1), and approximately the case where the video involves a live streamer which could potentially interacts with the viewers.
The time series $\{t_{j}^{e}\}$ of observed live chats of emotion $e$ is assumed to be sampled from a latent, inhomogeneous event intensity $\lambda^{e}(t)$. This intensity is influenced by three factors: (i) a homogeneous base-rate $\mu^{e}_{0}$ which captures live chat messages of emotion $e$ that are spontaneously posted by users in the absence of any influence from the video or live chat; (ii) an inhomogeneous rate $\mu^{e}_{1}(t)$ which captures live chat messages of emotion $e$ that are posted due to an influence from the video; (iii) finally, there is an endogenous rate accounting for the fact that past live chat messages trigger new ones. This is represented in the following expression for the intensity of emotion $e\in\mathcal{E}$ as a Hawkes point process:
@@
The intensity is defined such that $\lambda^{e}(t)dt$ is the probability that an emotion $e$ occurs in the live chat between $t$ and $t+dt$. Equation (1) states that the intensity at which emotions of type $e$ are generated in the live chat is an additive function of past emotions. The term $\mu_{0}^{e}$ is just a constant as described above. The term $S^{f}(t)$ is the time-varying presence of emotion $f$ in the video subscript (see Methods for details and Figure 1(d) for an example of $S^{\text{sad}}(t)$). The term $\nu^{e,f}S^{f}(t)$ then represents the rate at which emotion $f$ in the video triggers emotion $e$ in the live chat. For instance, a sad scene in the video can trigger an angry message in the live chat. Summing over all emotions $\mathcal{E}$ we arrive at the rate $\mu_{1}^{e}(t)$ by which any emotion from the video triggers emotion $e$ in the live chat. The third term captures the endogenous influence from prior live chat messages. Here, $\phi^{e,f}(t-t_{j}^{f})$ represents the influence of an emotion $f$ at time $t^{f}_{j}$ on emotion $e$ at time $t$. The intensity kernel $\phi^{e,f}(\cdot)$ is monotonously decaying, such that the more time has passed, the less of an influence a prior message has. We follow the common assumption that $\phi$ is exponential, that is $\phi^{e,f}(t)=\alpha^{e,f}\frac{1}{\gamma^{e}}e^{-\frac{1}{\gamma^{e}}t}$ with decay time $\gamma^{e}$. The larger $\gamma^{e}$ is, the longer the direct memory of a past live chat message has on emotion $e$. We assume that $\gamma^{e,f}\equiv\gamma^{e}$ for simplicity. The coefficient $\alpha^{e,f}$ represents the endogenous excitation effect from emotion $f$ to emotion $e$ in the live chat.

### v2

# Quantification of Interdependent Emotion Dynamics in Online Interactions

@@ Abstract
A growing share of human interactions now occurs online, where the expression and perception of emotions are often amplified and distorted. Yet, the interplay between different emotions and the extent to which they are driven by external stimuli or social feedback remains poorly understood. We calibrate a multivariate Hawkes self-exciting point process to model the temporal expression of six basic emotions in YouTube Live chats. This framework captures both temporal and cross-emotional dependencies while allowing us to disentangle the influence of video content (exogenous) from peer interactions (endogenous). We find that emotional expressions are up to four times more strongly driven by peer interaction than by video content. Positivity is more contagious, spreading three times more readily, whereas negativity is more memorable, lingering nearly twice as long. Moreover, we observe asymmetric cross-excitation, with negative emotions frequently triggering positive ones, a pattern consistent with trolling dynamics, but not the reverse. These findings highlight the central role of social interaction in shaping emotional dynamics online and the risks of emotional manipulation as human-chatbot interactions become increasingly realistic.
keywords:
non-normal system , transient dynamics , critical transitions , early-warning signals , tipping points
In recent decades, a growing share of human interaction has shifted to online platforms. The distinctive features of these environments, such as anonymity, reduced accountability, and the absence of non-verbal cues, shape digital interactions in ways that differ markedly from face-to-face communication [1]. This transformation has spurred substantial interest in understanding the role of emotions in online engagement and their broader societal consequences.
Emotional contagion in digital settings is well documented: Both positive and negative emotions propagate through social networks [2, 3, 4, 5, 6], and emotional content increases the likelihood of resharing [7, 3]. Moral emotions, in particular, enhance the spread of moral and political discourse [8], and the strength of social ties modulates emotional diffusion: strong ties amplify influence [9], while anger spreads more effectively than joy through weak ties, reaching wider audiences [10].
Platform design also shapes emotional dynamics. Digital media platforms are motivated to upregulate user emotions [11], and large-scale data from YouTube suggests that live-streaming environments can further intensify emotions through mechanisms such as shared attention [12]. Temporal studies show that positive emotions tend to rise quickly and fade fast, whereas negative emotions build more gradually and persist longer [13, 14, 15].
Emotional dynamics on social platforms also affect collective behavior and discourse by amplifying polarization, incivility, and misinformation [16, 17, 18, 19, 20, 21]. Negative emotions are particularly predictive of fake news engagement during crises, such as the COVID-19 pandemic [22]. Emotional expression in online reviews is also a good predictor of commercial outcomes across domains including movie and restaurant success [23].
Despite growing evidence of emotional contagion in digital spaces, we identify four key limitations in the literature. First, previous studies have found evidence for both positive and negative biases on online platforms, highlighting the need for further investigation to compare, contextualize, and reconcile these different effects [11, 24]. Online environments foster a positivity bias driven by self-representation and social validation mechanisms [25, 26, 27]. Meanwhile, negativity can spread more easily due to cognitive biases that prioritize negative information [28, 29, 30]. Previous studies have found that positivity tends to be more contagious on online platforms [4, 31, 32]. Meanwhile, recent findings suggest that while negativity often has greater reach [33, 34], emotional virality is complex and context-dependent [24]. These mixed results call for a reconciliation of our understanding of the different ways in which negative and positive emotions spread within social networks.
Second, while emotional contagion is well documented, less is known about how different emotions trigger one another in large-scale social interactions. Previous work has focused on individual emotion transitions [35, 36] and interpersonal dynamics, where emotional expressions can elicit both mimicry and divergent responses depending on the context [37, 38]. In addition, previous studies have used co-occurrence patterns of emotions to improve emotion classification using natural language processing (NLP) models [39, 40]. A graph-based approach incorporating emotion correlations has been found to outperform previous benchmarks in emotion classification tasks [41]. However, how different emotions trigger each other in the context of online social interactions is relatively underexplored.
Third, with few exceptions [13, 14, 15], most studies do not account for the subtle temporal dependencies underlying emotional contagion. Contagion is often measured in terms of whether content is shared or reciprocated, while the timing between emotional expressions receives comparatively little attention.
Fourth, the mechanisms by which online engagement shapes users’ emotional states remain inadequately understood. Although prior studies have shown that emotions are widely spread through social networks, it is still unclear whether this contagion is primarily content-driven arising from the emotional tone of the information, or peer-driven, shaped by social reinforcement and interactions among users. For example, live comments during real-time events exhibit greater emotional intensity compared to standard comments [12], but the extent to which this amplification stems from the emotional nature of the event versus social interactions among participants is unclear. More broadly, distinguishing between exogenous (e.g., content-driven) and endogenous (e.g., peer-driven) influences in online emotion contagion remains a fundamental challenge in the study of complex social systems [42].
@@
Figure 1: (a) Screenshot of a YouTube live video on the Theranos scandal involving Elizabeth Holmes with a live chat section on the right. The video content is captured by transcripts (subtitles), which we use to proxy exogenous emotional stimuli for the live chats. The live chat section displays timestamped messages from users reacting to the video content in real-time as the live video streams. We highlight an example of a transcript labeled as sad in blue, and a live chat message labeled as angry in red. (b,c) We visualize the extraction of emotions from the video and live chat in the screenshot above. We plot the arrival of live chat messages that are labeled as sad (angry), indicated with diamond markers in blue (red). The shared x-axis shows the time in the video in units of minutes. The y-axis is unit-less. We label emotions non-exclusively. For instance, the sentence “Holmes is a victim of the fake news media.” is labeled as both sad and angry. We assume these emotions are generated by the latent, inhomogeneous intensity defined by expression (1). (d) Time-varying component of the emotion sad in the video (transcript). Blue dots annotate the arrival of transcripts that are labeled as sad. The black dashed line represents the temporal function $S^{\text{sad}}(t)$, capturing the presence of sad emotions within the video.
We address these gaps by modeling emotional contagion as a sequence of discrete events—moments in time when users express distinct emotions in response to stimuli or peer interaction. Such event-based dynamics are naturally captured by self-exciting point processes, which allow past events to influence the probability of future ones. This temporal dependency and recursive structure make the multivariate Hawkes process a particularly suitable tool, as it explicitly accounts for feedback effects between events over time - a feature that standard regression approaches typically do not accommodate [43, 44].
Using this framework, we analyze the joint dynamics of six basic emotions in YouTube live chats. We classify emotions originating from video content as exogenous (exo) inputs and those expressed in the chat as endogenous (endo) responses driven by peer interaction. By estimating the parameters of a multivariate Hawkes model across 397 videos totaling more than 780 hours of content, we show that user emotional expressions are influenced by endogenous interactions roughly four times more than by exogenous video content. This suggests that in online environments, emotions are not simply reactions to external stimuli but are actively sustained and amplified through social feedback.
Our analysis further reveals a strong asymmetry in emotional contagion. Positive emotions, particularly joy, are about three times more likely to trigger additional joy than negative emotions are to trigger other negative expressions. However, this higher contagiousness is counterbalanced by slower temporal decay in negative emotions, which propagate over time windows nearly twice as long as those of positive ones. We also identify notable cross-excitation effects, especially cases where negative emotions trigger positive responses, a pattern reflective of antisocial dynamics such as trolling, where users derive gratification from provoking negative reactions [45, 46, 47].
Our study provides practical implications for the design of social networks and livestream platforms. First, increasing user exposure to joyful content can diffuse positivity across social communities. Second, active communication channels, such as live chats or bots, can much more effectively engage users emotionally compared to passive ones, such as video advertisements.
Methods
YouTube Live Chat Data
@@ Methods
We examine the dynamics of emotions in online discussions through a collection of YouTube Live videos. YouTube is a video sharing platform that caters to a wide range of users across different cultures. YouTube Live allows the audience to chat with each other while the video is being played in real time. We refer to live discussions as live chats and a single user live comment as a live chat message, or just a message. Videos with the live chat replay function record time-stamped live discussions and allow viewers to retrieve the live experience by simultaneously replaying the video content and live chat messages. Figure 1(a) shows an example of a YouTube live video (left) with a live chat section (right) where viewers post messages visible to the entire audience as the video streams.
We examine emotion dynamics in YouTube live videos with replay functionality across 27 topics ranging from comedy and documentary to sports and travel (see SI Appendix for the complete list). We first compile the list of topic keywords and collect completed live video IDs for each topic keyword using the keyword search function via the YouTube API in Python. 111 Code and data are available at https://github.com/ivylyis/Quantification-of-the-Self-Excited-Emotion-Dynamics-in-Online-Interactions. We initially extract both the video transcripts and live chat messages associated with each video. This process yields two distinct lists: one of time-stamped chat messages and another of time-stamped video transcripts (subtitles). We proceed by selecting only those videos in which the median interval between successive live chat messages ranges from 1 second to 5 minutes. Additionally, we disregard videos where more than 70% of the chat messages are in a language other than English. For the videos retained, we eliminate non-English live chat messages. Furthermore, we delete any chat messages that precede the first or succeed the last transcript time in each video. Ultimately, this results in a collection of 673,551 live chat messages from 1,957 videos.
Labeling Emotions in Text
@@
Multiple emotions can arise simultaneously in a single sentence [39, 41]. We thus assign to each text a vector of length 6 indicating the non-exclusive presence of each of the six emotions. This amounts to a multilabel classification problem. For training data, we rely on the SemEval-2018 dataset containing 6,838 Tweets with non-exclusive emotion labels across 11 emotions [53]. We group semantically related emotions into six core categories by mapping anticipation, optimism, love, and trust to joy; pessimism to sadness; and retaining sadness, anger, disgust, fear, and surprise in their original categories. Subsequently, we fine-tune a Roberta transformer model [54]. The transformer takes as input a text and gives as output the probabilities for each of the 6 basic emotions via 6 independent sigmoid activations. We use the binary cross-entropy loss function for training. During prediction, we convert each sigmoid output to a binary indicator at the 0.5 cutoff. This allows us to assign to each live chat message and video subscript a subset of the emotions $\mathcal{E}$ (see Figure 1 for an example).
For every video and emotion analyzed, we determine the average count of live chat messages per minute. To exclude outliers or inconsistencies in live chat data, only those videos where this average falls between the $20^{\text{th}}$- and $80^{\text{th}}$-quantiles of all videos are included. This process results in a compiled set of 92,412 live chat messages that span 397 videos, totaling 780 hours. The median video length is 107 minutes and per video, the median number of live chat messages is 204. When considering all videos, the median number of live chat messages per minute is as follows: 1.63 for joy, 0.17 for disgust, 0.16 for anger, 0.11 for sadness, and for both fear and surprise, 0.03. Additional details are provided in the SI Appendix.
Emotions as Self-Excited Point Processes
The treatment of emotions as discrete events makes point processes a natural analytical framework. A point process models the occurrence of discrete events in continuous time, with the simplest case being the Poisson process, where events occur independently at a constant rate. However, emotions are not independent; they exhibit self-excitation, which means that past emotions increase the likelihood of future ones [55]. To account for this, we adopt the Hawkes process, one of the simplest self-exciting point processes, which extends the Poisson process by incorporating memory and self-excitation while maintaining a linear and additive structure [56]. The intensity function of a Hawkes process is given by $\lambda(t)=\mu+\sum_{t_{i}<t}g(t-t_{i})$ where $\mu$ is the baseline event rate, and $g(t-t_{i})$ is a triggering function that quantifies the influence of past events that occurred at discrete times $\{t_{j}\}$ on future events. This formulation allows for a branching process interpretation, where events can be classified into exogenous (generated by external stimuli) and endogenous (triggered by past events) [57]. Hawkes processes have been widely applied to self-exciting phenomena, including earthquakes [58], financial transactions [59, 60], video viewing activities [61, 62], and information cascades on social networks [63, 64].
@@
Emotional contagion, the phenomenon in which emotions spread from one individual to another, is naturally captured by the self-exciting nature of the Hawkes process [55]. In our context, emotions are modeled as events whose statistical occurrence follows a point process. Because we have more than just one emotion and the emotions (presumably) cross-excite each other, we represent the dynamics of $|\mathcal{E}|=6$ distinct emotions using a multivariate Hawkes process [65]. It is our goal to understand to what extent a given emotion $f\in\mathcal{E}$ triggers another emotion $e\in\mathcal{E}$ either exogenously from the video to the live chat or endogenously from live chats to live chats. Here, the adjectives exogenous and endogenous refer to the point of view of live chat: If emotion $f$ appears in the video and triggers an emotion $e$ in the live chat, we call it exogenous (exo). In contrast, if emotion $f$ also appears in the live chat and subsequently triggers emotion $e$ in the live chat, we call it endogenous (endo). We assume that emotions in the video and in the live chat can trigger emotions in the live chat, but that emotions in the live chat cannot influence emotions in the video. This is precisely the case if the video is pre-recorded (e.g., a documentary, as in Figure 1), and approximately the case where the video involves a live streamer who could potentially interact with the viewers.
In a more formal representation of these concepts, we label $\{t_{j}^{e}\}$ as the set of timestamps corresponding to when a live chat message expressing emotion $e$ was sent. We assume that this set of event times is sampled from a latent inhomogeneous event intensity $\lambda^{e}(t)$. The intensity is determined by three factors: (i) a uniform, constant exo base rate $\mu^{e}_{0}$, reflecting the spontaneous generation of live chat messages of emotion $e$ without any prompting from the video or chat; (ii) a non-uniform, dynamic exo rate $\mu^{e}_{1}(t)$, representing the generation of live chat messages of emotion $e$ triggered by events in the video; (iii) an endogenous rate considers that earlier live chat messages provoke subsequent ones. This is defined by the following expression for the intensity of emotion $e\in\mathcal{E}$ as a Hawkes point process:
@@
Intensity is defined such that $\lambda^{e}(t)\mathrm{d}t$ is the probability that an emotion $e$ occurs in the live chat between $t$ and $t+\mathrm{d}t$. Equation (1) states that the intensity at which emotions of type $e$ are generated in the live chat is an additive function of past emotions. The term $\mu_{0}^{e}$ is just a constant as described above. The term $S^{f}(t)$ is the time-varying presence of emotion $f$ in the video subscript (see next section for details and Figure 1(d) for an example of $S^{\text{sad}}(t)$). The term $\nu^{e,f}S^{f}(t)$ then represents the rate at which emotion $f$ in the video triggers emotion $e$ in the live chat. For example, a sad scene in the video can trigger an angry message in the live chat. Summing over all emotions $\mathcal{E}$ we arrive at the rate $\mu_{1}^{e}(t)$. This rate $\mu_{1}^{e}(t)$ captures the rate at which any emotion in the video triggers emotion $e$ in the live chat. The third term captures the endogenous influence of previous live chat messages. Here, $\phi^{e,f}(t-t_{j}^{f})$ represents the influence of a past emotion $f$ at time $t^{f}_{j}$ on the likelihood of observing the emotion $e$ at time $t$. The intensity kernel $\phi^{e,f}(\cdot)$ is monotonically decaying, so the more time has passed, the less influence a prior message has. We follow the common assumption that $\phi$ is exponential, $\phi^{e,f}(t)=\left.\alpha^{e,f}e^{-t/\gamma^{e}}\right/\gamma^{e}$, with decay time $\gamma^{e}$. The larger $\gamma^{e}$, the longer the direct memory of a previous live chat message has on emotion $e$. Here, we have assumed $\gamma^{e,f}\equiv\gamma^{e}$ for simplicity, but generalizations are straightforward, albeit at the cost of increasing the number of parameters to be estimated. The coefficient $\alpha^{e,f}$ represents the endogenous excitation effect of emotion $f$ on emotion $e$ in the live chat.